• 4x2 - 12x + 9 =
  • 4(x2 - 3x + (9/4)) =
  • 4(x - 3/2)2 + 0


Kindly, tell me how this Result "4(x - 3/2)2 + 0" is obtained, please.


Thank you.



1
Expert's answer
2020-08-26T15:51:59-0400

1 step. 4x2 - 12x + 9 = 4(x2 - 3x + (9/4))  factoring out 4. To do it, divide each coefficient by 4.

2 step. 4*(x2 - 3x + (9/4)) = 4*(x - 3/2). According to the formula of full square: a2 - 2*a*b + b2 = (a - b)2

In your example 'a' is 1, 'b' is 3/2. Then a2 = 1, 2*a*b = 2*1*(3/2) = 3, b2 = (3/2)2 = 9/4

P.S. For every c : c+0 = c, so 0 can be cancelled in the answer.